OpenStudy (lcnettech):

Which options could be the next step in simplifying this expression (18 + 23) + 7 Choose exactly two answers that are correct. A. 41 + 7 (use order of operations) B. 25 + 30 (use the distributive property) C. 30 + 25 (use the commutative property) D. 18 + 30 (use the associative property)

OpenStudy (anonymous):

Unless I am wrong, all of them are. They all contain a variable, so I will say all of them.

OpenStudy (lcnettech):

i can only chose one

OpenStudy (anonymous):

C!

OpenStudy (anonymous):

I would also say c if you had to pick one.

OpenStudy (anonymous):

Here is what google says an algerbraic expression is, In mathematics, an algebraic expression is an expression built up from constants, variables, and the algebraic operations (addition, subtraction, multiplication, division and exponentiation by an exponent that is a rational number).

OpenStudy (lcnettech):

Which statement is true according to the order of operations? Choose exactly two answers that are correct. A. 6 + 10 ÷ 2 = 6 + 5 = 11 B. 8 – 4 ÷ 2 = 4 ÷ 2 =2 C. (11 + 3) • 2 = 14 • 2 = 28 D. 5 + (9 • 2) = 14 • 2 = 28

OpenStudy (lcnettech):

Which statement is true according to the order of operations? Choose exactly two answers that are correct. A. 6 + 10 ÷ 2 = 6 + 5 = 11 B. 8 – 4 ÷ 2 = 4 ÷ 2 = 2 C. (11 + 3) • 2 = 14 • 2 = 28 D. 5 + (9 • 2) = 14 • 2 = 28

OpenStudy (lcnettech):

the top one is the new question

OpenStudy (anonymous):

A and C are both right. PEMDAS-the order of operations. Parenthesis, Exponents, Multiplication and Division from left to right, Addition and Subtraction from left to right. Make sense?

OpenStudy (lcnettech):

yep.can you help with a few more?

OpenStudy (anonymous):

If I know how to do them! Then yes.

OpenStudy (anonymous):

I have until 10:55 a.m. (eastern time)

OpenStudy (lcnettech):

(5 * 8 - 12)(36 divided by 9 * 4)

OpenStudy (lcnettech):

what time is it right now?

OpenStudy (anonymous):

448 i believe. It is 10:50.

OpenStudy (lcnettech):

k

OpenStudy (anonymous):

I can maybe answer one more problem.

OpenStudy (anonymous):

I do not mean to be rude, but I have to go. Sorry.

OpenStudy (lcnettech):

thats ok